### v1

@@ 5.1.2 Prompt using External Knowledge
We can use external information about the problem to improve the prompt. Press et al. (2022) study how subproblems are related to the main problem, which they call compositional reasoning. They study how often a model is able to answer the subproblems, but not the overall problem. This difference is called the compositionality gap. They find that in GPT-3, as model size increases, the compositionality gap does not decrease: the single-hop question-answering performance improves faster than the multi-hop performance. This shows that while more powerful models memorize and recall more factual knowledge, no improvement in their ability to perform compositional reasoning occurs. They find that the ability to reason does not depend on the size of the model.
@@
Subsequently, a method called Self-ask is proposed, that asks elicitive follow-up questions (like Chain-of-thought, but with the follow up: prompt), see Figure 5. The model is then used to answer these follow-up questions. Self-ask can also use an external search engine to answer intermediate prompts, instead of the model. The model takes as input a compositional question which it decomposes. The initial subquestion is fed into the search engine, and the answer is processed by the model, which generates another subquestion, and so on, until it produces the final answer.
The approach performs a few percentage points better than vanilla Chain-of-thought on three benchmarks that were specifically designed for multi-hop questions.
5.1.3 Model-Generated Prompt
@@
In addition to manually writing prompts or using external information, we can also try to let the LLM itself study the problem to write the best reasoning-prompt, a form of self-improvement. An example of this approach is Auto-chain-of-thought (Zhang et al., 2022). This approach builds on the observation by Kojima et al. (2022) that large language models are zero-shot reasoners. First, Auto-chain generates specific questions for a given dataset and partitions them into clusters. Then an external algorithm uses the model to generate examples that are sampled for diversity. The constructed demonstrations augment the in-context prompt. The automatically generated prompts are reported to perform as well or better than the hand-written Chain-of-thought prompts on ten benchmarks using GPT-3.
@@
Fu et al. (2022) introduce Complexity-based prompting. Inspired by Chain-of-thought and Self-consistency, this work studies which prompts achieve the best results on math word and other reasoning problems. Their work specifically studies the impact of the complexity of the reasoning chain, and introduces a related reasoning approach (Complexity-based prompting). They find that prompts with the largest complexity (the most reasoning steps) perform best. Further, they find that outputs (answers) with the highest complexity are the best. Complexity-based prompting achieves high performance on three math reasoning benchmarks.
Another approach that uses model-generated prompts is Buffer-of-thoughts. We will discuss this approach in Section 5.3.3.
5.2 Evaluation of Steps
@@
After discussing prompts for the generation of reasoning steps, the next stage in the reasoning pipeline (Section 2.3) is evaluation of the results of the steps, to reduce the error of multi-step reasoning chains.
We will start with approaches where the same model performs step-generation and step-evaluation.
5.2.1 Self-Assessment
When LLMs are prompted to perform reasoning steps, they perform a sequence of steps and predict multiple tokens. Performing a sequence of steps makes them sensitive to mistakes and vulnerable to error accumulation (Weng et al., 2022; Xiao et al., 2023a). Several methods have been developed to prevent error accumulation. One approach is to create a new model to separately evaluate the results. Shen et al. (2021) and Li et al. (2022b) train an external verifier to check results.
@@ 5.2.1 Self-Assessment
In contrast, Weng et al. (2022) propose an automated approach using evaluation by the same LLM, called Self-verification. They note that human reasoning also suffers from the problem of accumulating errors, and that in human reasoning we frequently revisit our thought process to verify the accuracy of our reasoning steps. Thus, they propose to apply such a backwards self-verification approach. The LLM is prompted to use the conclusion of the Chain-of-thought reasoning chain as a condition for solving the original problem and then compare the answer going back to the original question. The LLM is given variations of its own conclusion and is instructed to choose the one with the highest similarity to the original question. (Note that there can be feedback issues using an LLM to evaluate itself, for a discussion see Zheng et al. (2024).) Experiments are reported on GPT-3 (Chen et al., 2021) and on Instruct-GPT (Ouyang et al., 2022). The performance of Chain-of-thought was improved by a few percentage points on arithmetic and general reasoning tasks.
@@
A popular related approach is called Self-consistency (Wang et al., 2022b). Self-consistency is a straightforward ensemble approach. Greedy single-path decoding is replaced by sampling diverse reasoning paths, evaluating them, and selecting the most consistent answer. Self-consistency asks the LLM to simply perform the same query multiple times, and takes the majority-vote of the answers. Self-consistency works since complex reasoning problems typically allow different reasoning paths that lead to the correct answer. Figure 6 summarizes the approach.
@@
Self-consistency has been evaluated on arithmetic reasoning, commonsense reasoning and symbolic reasoning, on a variety of LLMs, including GPT-3 (Tay et al., 2022; Brown et al., 2020; Thoppilan et al., 2022; Chowdhery et al., 2023). Self-consistency improves the performance of Chain-of-thought typically by 10-20 percentage points, and has been used as a baseline in many of the other approaches in this survey. (Self-verification also reports that performance is improved when used in combination with Self-consistency (Wang et al., 2022b) and with Program-aided-language (Gao et al., 2023).)
5.2.2 Tool-based Validation
@@
Another possibility to improve the accuracy of evaluating the reasoning steps is to switch from a natural to a formal language. The advantage of a formal language is that it is less ambiguous than a natural language. Examples are computer languages, such as Python, or mathematical equations. Using a formal language for reasoning is a popular approach, and we discuss seven papers. Many approaches generate the steps in Python, and the code can then be evaluated by a formal evaluator, such as a compiler, debugger, or interpreter.
@@
LLMs have been quite successful in generating computer code from natural language prompts. Chen et al. (2021) introduced Codex, a GPT model that was trained on publicly available code in the repository GitHub. A production version of this work was introduced under the name GitHub Copilot. Codex is able to generate correct programs from descriptions in natural language, such as commentary strings. Figure 7 shows examples that are produced by Codex.
The work on Codex is used as a basis for further research on reasoning in LLMs.
@@
Human programmers, when writing code, typically follow a cycle of writing some code, executing it to look for errors, and then using the feedback to improve the code. This same approach is followed in the Self-debugging work (Chen et al., 2023). Self-debugging teaches a large language model to debug its generated program code via few-shot demonstrations. It follows the same steps of (1) code generation, (2) code execution, and (3) code explanation (see Figure 8).
Self-debugging is able, without human feedback on the code’s correctness or error messages, to identify mistakes in the code that was generated by itself from investigating the execution results. Self-debugging can also provide an explanation of the generated code in natural language. It achieves strong performance on text-to-SQL generation, C++-to-Python transcoding, and text-to-Python generation.
@@
Several works use self-debugging to generate working code tuned for solving specific problems automatically, without human feedback. Romera-Paredes et al. (2024) introduced FunSearch, an approach that integrates formal methods and LLMs to enhance mathematical reasoning and code generation. FunSearch is capable of producing functionally correct programs that adhere to specified requirements. It uses a genetic algorithm approach with multiple populations of candidate solutions (programs), which are automatically evaluated (using tools depending on the problem specification). In addition to the problem specification in the form of an evaluate function, also an initial program is given to the LLM in the first prompt. After evaluating a number of generated programs from the starting prompt, a new prompt using ‘best-shot prompting’ is created in an iterative fashion, combining a selection of sampled programs in a sorted list (ascending according to their evaluation score), and the LLM is requested to generate program . Another work leverages evolutionary computation methods to generate and optimize evolutionary algorithms (van Stein and Bäck, 2024). This approach, LLaMEA (Large Language Model Evolutionary Algorithm), utilizes LLMs to design and optimize evolutionary algorithms. The approach uses LLMs to generate initial algorithmic structures, which are then refined through mutation and selection. This enhances the efficiency of algorithm design, particularly in fields requiring innovative and adaptive solutions. A key difference between FunSearch and LLaMEA is that LLaMEA uses a sample-efficient elitism strategy by iteratively improving the best-so-far solution, requiring significantly fewer prompt evaluations than the large-population strategy proposed in FunSearch.
@@
To improve prompt-based reasoning, Codex is used in an ensemble approach named MathPrompter (Imani et al., 2023). This approach generates multiple algebraic expressions or Python functions, which then solve the same math problem. The results are compared, just like in Self-consistency and Self-verification, raising the confidence level in the results. MathPrompter achieved state-of-the-art results on the MultiArith dataset (78.7% 92.5%), evaluated on GPT-3 175B.
Two other approaches that use a formal language are Program-of-thought (PoT) (Chen et al., 2022) and Program-aided-language (PAL) (Gao et al., 2023). Both approaches use the LLM to generate Python and then use a Python interpreter to evaluate the result. PoT and PAL are similar approaches. PoT uses benchmark-specific prompts; PAL uses generic prompts, and has been tested on more benchmarks and has been used in other approaches. Figure 9 illustrates the PAL approach.
@@
When the evaluation of the reasoning steps is offloaded to the Python interpreter, decomposing the natural language problem into executable code-steps remains the only task for the LLM. (Earlier work in mathematical word problems, such as Ling et al. (2017), showed how to decompose a problem and reach an answer.) Gao et al. (2023) provide extensive experimental evidence about the synergy between the neural LLM and the symbolic interpreter. Experiments are performed over 13 mathematical, symbolic, and algorithmic reasoning tasks, achieving more accurate results than much larger models.
5.2.3 External Model Validation
@@
We have seen many successful examples of prompt-based in-context reasoning and evaluation. We will now look at related reasoning approaches that follow a more traditional parameter learning approach. We describe three natural language approaches that follow this route. All approaches evaluate the output of the model and generate corrective data. That data is then added to the training pipeline, and the model is subsequently finetuned.
@@
The Refiner approach (Paul et al., 2023) uses a generator model and a critic model to provide fine-grained feedback on reasoning errors. The generator generates multiple reasoning hypotheses, the critic evaluates results by randomly selecting a hypothesis for feedback. The generator model is finetuned based on its reasoning errors. A small supervised model is used to overcome the cold-start problem. Figure 10 shows an example of how the critic provides feedback to the generator.
The approach is reported to work well on math word problems and synthetic natural language reasoning.
Welleck et al. (2022) follow a similar approach in their Self-correction approach. The corrector is a separate model specialized in refining the outputs of the generator. Unlike Refiner, where the generator is finetuned based on the critic feedback, Self-correction finetunes the corrector to rectify errors in the hypotheses produced by the generator.
A third finetuning approach is Self-improvement, by Huang et al. (2022a). Here too the base model data is augmented by LLM-generated rationales, and then finetuned. Noteworthy in all three finetuning approaches is that LLMs are capable of improving themselves by training on their own generated output, and that stability problems inherent in feedback loops are overcome.
@@
The final finetuning approach that we discuss uses dataset augmentation. An explicit intermediate reasoning is called a rationale. Rationale generation has been shown to be valuable for LLMs across diverse tasks such as mathematical and commonsense reasoning, code evaluation, social bias inference, and natural language inference (Zelikman et al., 2022). Zelikman et al. (2022) describe how reasoning steps are used to create rationales, that are then used to augment the dataset on which the model is finetuned. The approach is called Self-taught-reasoner. Figure 11 illustrates the approach.
In Self-taught-reasoner, an augmentation dataset is created by attempting to solve the original dataset using the current model’s rationale generation ability in each iteration. Next, the dataset is augmented using rationalizations, using ground-truth answers to problems the model failed to solve. Finally, the large language model is finetuned on the combined dataset.
@@
In addition to math word problems, prompt-based reasoning has also been used to reason about robot behavior. Language models contain a large amount of information about the real world (Ahn et al., 2022). In theory, this should allow the model to exhibit realistic reasoning about robotic behavior. However, the models do not have knowledge about particular embodied aspects of a particular robot. If we could compare a Scratchpad-like list of intermediate reasoning steps with a list of possible movements of the robot in its environment, then we could prevent the model from suggesting impossible joint movements and actions, and prevent accidents.
Such an approach has been tried in the Say-can paper (Ahn et al., 2022). Say-can learns a value function (Kaelbling et al., 1996) of the behavior of a robot in an environment using temporal difference reinforcement learning Sutton (1988). This value function is then combined with prompt-based reasoning by the language model, to constrain it from suggesting impossible or harmful actions.
The goal of Say-can is to ground language in robotic affordances. In contrast to Scratchpad, which used supervised learning, the affordance model is learned interactively by reinforcement learning, and then applied using prompt-based learning by the LLM. The robot can act as the language model’s hands and eyes, while the language model has high-level semantic knowledge about the task. The LLM (Say) provides a task-grounding to find the actions to achieve the high-level goal. The learned affordance function (Can) provides a world-grounding to allow what is possible. Say-can is evaluated on 101 real-world robotic tasks, such as how to solve tasks in a kitchen environment (see Figure 12).
Where Say-can learns affordance as a separate function, another approach, Inner-monologue (Huang et al., 2022b) formulates robotic planning directly as part of the language prompt. This approach incorporates environmental information into the prompt, linguistically, as an inner monologue. As in Say-can, the information comes as feedback from different sources. Unlike Say-can, the information of physics and the world is inserted directly into the prompt.
Inner-monologue consists of many elements: it uses InstructGPT (Brown et al., 2020) for multi-step planning, scripted modules for object recognition, success detection, task-progress scene description, and language-conditioned pick-and-place primitives, similar to CLIPort (Shridhar et al., 2022). These elements generate textual descriptions that are used in prompt-based learning. Figure 13 gives an example of the working of Inner-monologue.
The language feedback that is thus generated significantly improves performance on three domains, such as simulated and real table top rearrangement tasks and manipulation tasks in a kitchen environment. There are many studies into robotic behavior. A recent approach related to Inner-monologue is Chain-of-tools, which proposes a plan-execute-observe pipeline to ground reasoning about tool behavior (Shi et al., 2024a, b).
@@
5.3 Control of Steps
The third stage in the reasoning pipeline in Section 2.3 is reasoning control. This stage controls how many sub-steps are generated, and how deep into the future the reasoning chain is generated.

### v2

@@ 3.1.2. Prompt using External Knowledge
Chain-of-thought prompts are written manually, by the researchers, an approach that does not scale. We can also use external information about the problem to improve the prompt. Press et al. (2023) study how subproblems are related to the main problem, which they call compositional reasoning. They study how often a model is able to answer the subproblems, but not the overall problem. This difference is called the compositionality gap. They find that in GPT-3, as model size increases, the single-hop question-answering performance improves faster than the multi-hop performance: while more powerful models memorize and recall more factual knowledge, no improvement in their compositional reasoning occurs. The ability to reason does not depend on the size of the model.
@@
Subsequently, a method called Self-ask is proposed, that asks elicitive follow-up questions (like Chain-of-thought, but with the follow up: prompt), that the model then answers. Self-ask can also use an external search engine to answer intermediate prompts, instead of the model. The initial subquestion is fed into the search engine, and the answer is processed by the model, which generates another subquestion, and so on, until it produces the final answer. Self-ask was tested on three benchmarks that were specifically designed for multi-hop questions. Although it performs only a few percentage points better than vanilla Chain-of-thought, it showed how external knowledge can be used in a reasoning setting.
3.1.3. Model-Generated Prompt
@@
In addition to manually writing prompts or using external information, we can also let the LLM itself study the problem to write the best reasoning-prompt. An example of such self-improvement is Auto-chain-of-thought (Zhang et al., 2023). This approach builds on the observation by Kojima et al. (2022) that large language models are zero-shot reasoners. First, Auto-chain generates specific questions for a given dataset and partitions them into clusters. Then an external algorithm uses the model to generate examples that are sampled for diversity. The constructed demonstrations augment the in-context prompt. This approach also performed a few percentage points better than hand-written Chain-of-thought prompts, on ten benchmarks, using GPT-3 (see Table 2).
@@
Fu et al. (2022) introduce Complexity-based prompting. Inspired by Chain-of-thought and Self-consistency, their work specifically studies the impact of the complexity of the reasoning chain (the number of steps), and introduces a related reasoning approach (Complexity-based prompting). They find that prompts with the largest complexity perform best, and also that answers with the highest complexity are the best. Complexity-based prompting achieves somewhat higher performance on three math reasoning benchmarks: GSM8K improves 7 points, MathQA 6 points, and the Penguins benchmark from Big Bench Hard improve 3 percentage points.
We see that the initial approaches showed larger improvements than the later approaches. It is time to look at another category of approaches, that focus on the evaluation of reasoning steps.
3.2. Evaluation of Steps
@@
After discussing prompts for the generation of reasoning steps, the next stage in the generation/evaluation/control pipeline is evaluation of the results of the steps. This stage focuses on reducing error accumulation of multi-step reasoning chains. We will start with approaches where the same model performs step-generation and step-evaluation.
3.2.1. Self-Assessment
@@
When LLMs are prompted to perform reasoning steps, they perform a sequence of steps and predict multiple tokens. Performing a sequence of steps makes them sensitive to mistakes and vulnerable to error accumulation (logical, factual, ethical, or otherwise) (Weng et al., 2023; Xiao et al., 2023a). Several methods have been developed to prevent error accumulation. One approach is to create a new model to separately evaluate the results. Shen et al. (2021) and Li et al. (2023b) train an external verifier to check results. In contrast, Weng et al. (2023) propose an automated approach using evaluation by the same LLM, called Self-verification. They note that human reasoning also suffers from the problem of accumulating errors, and that in human reasoning we frequently revisit our thought process to verify the accuracy of our reasoning steps. The LLM is prompted to use the conclusion of the Chain-of-thought reasoning chain as a condition for solving the original problem and then compare the answer, going back to the original question. The LLM is given variations of its own conclusion and is instructed to choose the one with the highest similarity to the original question. (Note that there can be feedback issues using an LLM to evaluate itself, for a discussion see Zheng et al. (2024).) Experiments are reported on GPT-3 (Chen et al., 2021) and on Instruct-GPT (Ouyang et al., 2022). The accuracy of Chain-of-thought was improved by a few percentage points on arithmetic and general reasoning tasks (GSM8K 65%, AQuA 48%, SVAMP 77%).
@@
A popular related approach is Self-consistency (Wang et al., 2023b). Self-consistency is a straightforward ensemble approach (a well-known machine learning technique to make a strong learner out of multiple weaker learners (Sagi and Rokach, 2018; Breiman, 2001)). Greedy single-path decoding is replaced by sampling diverse reasoning paths, evaluating them, and selecting the most consistent answer. Self-consistency asks the LLM to simply perform the same query multiple times, and takes the majority-vote of the answers, or decoding paths. Self-consistency works since complex reasoning problems typically allow different reasoning paths that lead to the correct answer. Figure 4 summarizes the approach.
@@
Self-consistency has been evaluated on arithmetic reasoning, commonsense reasoning and symbolic reasoning, on a variety of LLMs, including GPT-3 (Tay et al., 2023; Brown et al., 2020; Thoppilan et al., 2022; Chowdhery et al., 2023). Self-consistency further improves the performance of Chain-of-thought by 10-20 percentage points, and has been used as a baseline in many of the other approaches in this survey.
3.2.2. Tool-based Validation
@@
Another approach to improve the accuracy of evaluating the reasoning steps is to switch from a natural to a formal language. The advantage of a formal language is that it is less ambiguous than a natural language. Examples are computer languages, such as Python, or mathematical equations. Using a formal language for reasoning is a popular approach, and we discuss seven papers. Many approaches generate the steps in Python, and the code can then be evaluated by a formal evaluator, such as a compiler, debugger, or interpreter.
@@
LLMs have been quite successful in generating computer code from natural language prompts. Chen et al. (2021) introduced Codex, a GPT model that was trained on publicly available code in the repository GitHub. A production version of this work was introduced under the name GitHub Copilot. Codex is able to generate correct programs from descriptions in natural language, such as commentary strings.
@@
The work on Codex is used as a basis for further research on reasoning in LLMs. Human programmers, when writing code, typically follow a cycle of writing some code, executing it to look for errors, and then using the feedback to improve the code. This step-by-step approach is followed in Self-debugging (Chen et al., 2024). It follows the same steps of code generation, code execution, and code explanation. Self-debugging is able to identify mistakes in its own code by investigating the execution results, and can also provide an explanation of the generated code, in natural language. It achieves strong performance: the text-to-SQL Spider benchmark improves 9 points, and the C++ to Python Transcoder benchmark improves by 12 percentage points.
@@
Several works generate working code tuned for solving specific problems automatically, without human feedback. Romera-Paredes et al. (2024) introduced FunSearch, an approach that integrates formal methods and LLMs to enhance mathematical reasoning and code generation. FunSearch uses a genetic approach with multiple populations of candidate solutions (programs), which are evaluated using a function depending on the problem specification. In addition to the evaluation function, also an initial program is given to the LLM in the first prompt. After evaluating a number of generated programs from the starting prompt, a new prompt is created, in an iterative fashion, combining a selection of sampled programs sorted according to their evaluation score, and the LLM is requested to generate a new program. Another work leverages evolutionary computation methods to generate and optimize evolutionary algorithms (van Stein and Bäck, 2024). This approach, LLaMEA (Large Language Model Evolutionary Algorithm), utilizes LLMs to design and optimize evolutionary algorithms. The approach uses LLMs to generate initial algorithmic structures, which are then refined through mutation and selection. This enhances the efficiency of algorithm design, particularly in fields requiring innovative and adaptive solutions, improving accuracy on the Black-Box Optimization Benchmark (Hansen and Ros, 2010) (BBOB) suite by 11 pecentage points. A key difference between FunSearch and LLaMEA is that LLaMEA uses a sample-efficient elitism strategy by iteratively improving the best-so-far solution, requiring significantly fewer prompt evaluations than the large-population strategy proposed in FunSearch. Evolutionary approaches let the LLM discover new algorithms, solving existing problems in new ways, or solving entirely new problems. Another method, Evolution-of-heuristics (Liu et al., 2024), was proposed for evolving code snippets for guided local search to solve combinatorial optimization problems, such as the Traveling Salesperson Problem.
@@
To improve prompt-based reasoning, Codex is used in an ensemble approach named MathPrompter (Imani et al., 2023). This approach generates multiple algebraic expressions or Python functions, which then solve the same math problem. The results are compared, just like in Self-consistency and Self-verification, raising the confidence level in the results. MathPrompter achieved state-of-the-art accuracy on the MultiArith dataset (78.7% 92.5%), evaluated on GPT-3 175B.
@@
Two other approaches that use a formal language are Program-of-thought (Chen et al., 2023) and Program-aided-language (Gao et al., 2023). Both approaches use the LLM to generate Python and then use an interpreter to evaluate the result. The approaches are similar although Program-aided-language uses generic prompts, and has been tested on more benchmarks. Figure 3 (fourth column) illustrates the Program-aided-language approach. When the evaluation of the reasoning steps is offloaded to the Python interpreter, decomposing the natural language problem into executable code-steps remains the only task for the LLM. (Earlier work in math word problems showed how to decompose a problem and reach an answer (Ling et al., 2017).) Gao et al. (2023) provide extensive experimental evidence about the synergy between the neural LLM and the symbolic interpreter. Experiments are performed on 13 mathematical, symbolic, and algorithmic reasoning tasks, achieving more accurate results than much larger models (for example, Program-aided-language reported 72% on GSM8K and 93% on Penguins).
3.2.3. External Model Validation
@@
We have seen many examples of succesful prompt-based in-context reasoning and evaluation (at inference time—where no parameters were changed). We will now look at reasoning approaches that follow a more traditional parameter training approach. All approaches evaluate the output of the model and generate corrective data. That data is then added to the training pipeline, and the model is subsequently finetuned.
@@
The Refiner approach (Paul et al., 2024a) uses a generator model and a critic model that provide fine-grained feedback on reasoning errors. The generator generates multiple reasoning hypotheses, and the critic evaluates results by randomly selecting a hypothesis for feedback. The generator model is then finetuned based on its reasoning errors. A small supervised model is used to overcome the cold-start problem. The approach achieves 78% accuracy on GSM8K and also works well on related problems.
Welleck et al. (2023) follow a similar approach which they call Self-correction. Here, the corrector is a separate model specialized in refining the outputs of the generator. Unlike Refiner, where the generator is finetuned based on the critic, Self-correction finetunes the corrector to rectify errors in the hypotheses produced by the generator. Self-corrector is not applied to math word problems, but to program synthesis, where a small corrector reduces toxicity to 0%.
A third finetuning approach is Self-improvement (Huang et al., 2023). Here, too, the base model data is augmented by LLM-generated rationales, and then finetuned. They achieve 82% accuracy on GSM8K and similarly high scores on question answering and adversarial benchmarks. Noteworthy in all three finetuning approaches is that LLMs are capable of improving themselves by training on their own generated output, and that stability problems from feedback loops are overcome.
@@
The final finetuning approach that we discuss uses dataset augmentation. In Self-taught-reasoner (Zelikman et al., 2022), an intermediate reasoning is generated, called a rationale. Rationales are shown to be valuable across diverse tasks such as mathematical and commonsense reasoning, code evaluation, social bias inference, and natural language inference. First an augmentation dataset is created by attempting to solve the original dataset. Next, the dataset is augmented using rationalizations and ground-truth answers to problems the model failed to solve. Finally, the model is finetuned on the combined dataset. Self-taught-reasoner performs comparably (72%) to finetuning a 30 times larger model on CommonsenseQA.
@@
In addition to math word problems, computer code, and common sense, prompt-based reasoning has also been used to improve robot behavior. Language models contain a large amount of information about the real world (Ahn et al., 2022). In theory, this should allow them to reason realistically about robotic behavior. However, the models do not have knowledge about specific embodied aspects of a particular robot. If we could compare a Scratchpad-like list of intermediate reasoning steps with a list of possible movements of the robot in its environment, then we could prevent the model from suggesting impossible joint movements and actions, and prevent failures.
Say-can (Ahn et al., 2022) learns a value function (Kaelbling et al., 1996) of the behavior of a robot and its environment using temporal difference reinforcement learning (Sutton, 1988). This value function is combined with prompt-based reasoning by the language model, to constrain it from suggesting impossible actions. The goal of Say-can is to ground language in robotic affordances. In contrast to Scratchpad, which used supervised learning, the affordance model is learned interactively by reinforcement learning, and then applied using prompt-based learning on the LLM. The language model has high-level semantic knowledge about the task (Say). The learned affordance function (Can) provides an environment-grounding on what is possible. Say-can achieves a 31% success rate on 101 real-world robotic kitchen tasks.
Where Say-can learns affordances as a separate function, Inner-monologue (Huang et al., 2022) formulates robotic planning directly as part of the language prompt, internally. The input consists of many elements: textual descriptions from InstructGPT (Brown et al., 2020) for multi-step planning, scripted modules for object recognition, success detection, task-progress scene description, and language-conditioned pick-and-place primitives, similar to CLIPort (Shridhar et al., 2022). The language feedback that is thus generated significantly improves performance on three benchmarks, achieving 90% accuracy on simulated and real table top rearrangement tasks and 60% on the kitchen environment. There are many other studies into robotic behavior. An approach related to Inner-monologue is Chain-of-tools, which proposes a plan-execute-observe pipeline to ground reasoning about tool behavior (Shi et al., 2024a, b).
@@
3.3. Control of Steps